### v2

# Deep Generative Models in Condition and Structural Health Monitoring: Opportunities, Limitations and Future Outlook

Abstract
Condition and structural health monitoring (CM/SHM) is a pivotal component of predictive maintenance (PdM) strategies across diverse industrial sectors, including mechanical rotating machinery, aircraft structures, wind turbines, and civil infrastructures. Conventional deep learning models, while effective for fault diagnosis and anomaly detection through automatic feature learning from sensor data, often struggle with operational variability, imbalanced or scarce fault datasets, and multimodal sensory data from complex systems. Deep generative models (DGMs) including deep autoregressive models, variational autoencoders, generative adversarial networks, diffusion-based models, and emerging large language models, offer transformative capabilities by synthesizing high-fidelity data samples, reconstructing latent system states, and modeling complex multimodal data streams. This review systematically examines state-of-the-art DGM applications in CM/SHM across the four main industrial systems mentioned above, emphasizing their roles in addressing key challenges: data generation, domain adaptation and generalization, multimodal data fusion, and downstream fault diagnosis and anomaly detection tasks, with rigorous comparison among signal processing, conventional machine learning or deep learning models, and DGMs. Lastly, we discuss current limitations of DGMs, including challenges of explainable and trustworthy models, computational inefficiencies for edge deployment, and the need for parameter-efficient fine-tuning strategies. Future research directions can focus on zero-shot and few-shot learning, robust multimodal data generation, hybrid architectures integrating DGMs with physics knowledge, and reinforcement learning with DGMs to enhance robustness and accuracy in industrial scenarios.
keywords:
Predictive maintenance, Condition monitoring, Structural health monitoring, Deep generative models, Large language models, Diffusion modelsNomenclature
- Acronyms
CM
Condition monitoring
DA
Domain adaptation
DAR
Deep autoregressive
DDIMs
Denoising diffusion implicit models
DDPMs
Denoising diffusion probabilistic models
DGMs
Deep generative model
DNNs
Deep neural networks
ELBO
Evidence Lower Bound
FD&AD
Fault diagnosis and anomaly detection
FID
Fréchet inception distance
GAN
Generative adversarial network
GenAI
Generative artificial intelligence
GenXAI
Explainable Generative AI
GMMs
Gaussian mixture models
GPT-4
Generative pre-trained transformer 4
GRU
Gated recurrent unit
HMMs
hidden Markov models
KNN
K-nearest neighbors
LLM
Large language model
LoRA
Low-Rank Adaptation
LSTM
Long short-term memory
NDT
Non-destructive testing
NLL
Negative log-likelihood
PdM
Predictive maintenance
PEFT
Parameter-efficient fine-tuning
PINNs
Physics-informed neural networks
PTQ
Post-training quantization
QLoRA
Quantized LoRA
RAG
Retrieval-augmented generation
RL
Reinforcement learning
RNN
Recurrent neural networks
SCADA
Supervisory Control and Data Acquisition
SHAP
SHapley Additive exPlanations
SHM
Structural health monitoring
SMOTE
Synthetic minority over-sampling technique
SVR
Support vector regression
TL
Transfer learning
VAE
Variational autoencoder
- List of Symbols
Noise scale
Weight update matrix
Expectation
The set of real numbers
Target sample
Latent vector
Constant
The identity matrix
Key matrices
Query matrices
Value matrices
Weight matrices
Kullback–Leibler divergence
Mapping generator
Loss function
Gaussian distribution
Target distribution
Tractable distribution
Training dataset
Mean value
Covariance matrix
Model parameters
Gaussian noise variable
Amplitude
GAN discriminator
GAN generator
Hidden state
Number of data samples
Probability distribution
Backward transition probability
Forward transition probability
Time steps
Contents
1 Introduction
1.1 Background
Safety and reliability are of great importance for industrial operational systems, driving the need for predictive strategies to mitigate risks and optimize performance. Predictive maintenance (PdM) has emerged as a pivotal strategy, leveraging intelligent data analysis and monitoring techniques to prevent potential failures, minimize downtime, and prolong equipment lifespans [1, 2, 3]. As a core component of PdM, condition monitoring (CM), also termed condition-based maintenance [4], is a real-time process involving sensor-based technologies to track equipment parameters such as vibration, temperature and acoustics in order to assess system health and detect early anomalies. Therefore, CM also involves fault diagnosis and anomaly detection (FD&AD) tasks using real-time data and enables immediate corrective actions to prevent failures. Specifically, in rotating machinery systems, CM stands out by identifying subtle deviations from normal operating profiles, detecting incipient faults (e.g., bearing wear, gear degradation, misalignment due to improper installation, asymmetric loads). These insights are pivotal for rotating systems, where unexpected failures can lead to safety hazards, production losses, or environmental impacts. For static structures, such as pipelines, bridges, or fixed machinery, structural health monitoring (SHM) and non-destructive testing (NDT) are frequently used to assess structural integrity without disrupting operations [5, 6]. NDT techniques, including ultrasonic testing and thermography, assess structural safety without compromising functionality by identifying defects such as cracks or corrosion. SHM expands on this by deploying permanent sensor networks to continuously monitor static systems for gradual degradation.
For intelligent CM/SHM systems, deep neural networks (DNNs) are widely utilized for FD&AD tasks. Representative frameworks, such as convolutional neural networks [7], recurrent neural networks [8, 9], and more recent Transformer-based networks [10], have demonstrated effectiveness in FD&AD tasks by extracting useful features from heterogeneous data modalities, including images, acoustic signals, and time series sensor data. However, limitations and challenges still exist for applying DNNs within industrial scenarios:
Scarce or missing data: Prevalent DNNs are limited by their dependence on large, high-quality labeled datasets to achieve robust classification or prediction performance. Industrial environments frequently struggle to collect sufficient labeled data, or the presence of missing data due to sensor failure, communication interruption, and environmental interference [11, 12];
Domain mismatch: Sensor data collected under normal and abnormal conditions often exhibits severe class imbalance, as industrial systems predominantly operate within nominal states during routine cycles. Even when fault data are available, they may display uneven distribution across distinct fault categories, compromising model training and generalization [13].
Multimodal data complexity: Industrial sensor data originate from heterogeneous sources, such as vibration time series, thermal imaging, acoustic signals, and emission sensors. These multimodal datasets require advanced techniques to address challenges in feature alignment, temporal synchronization, and cross-modal fusion [12].
As a promising solution, more recently, deep generative models, under the umbrella of generative artificial intelligence (GenAI), have attracted broad interest in addressing the challenges in the CM/SHM field.
1.2 Deep generative models
Deep generative models (DGMs) are deep neural networks with multiple hidden layers trained to approximate complicated, high-dimensional probability distributions of training samples [16, 17]. These models have become the dominant paradigm in GenAI, driven by advances in stochastic optimization and the availability of massive training datasets [18, 19, 20]. The term deep indicates the use of deep neural networks as parametric models in the approximation process [21]. To anchor the view in industrial health monitoring, this review focuses on five prominent DGM frameworks: deep autoregressive (DAR) models, variational autoencoders (VAEs), generative adversarial networks (GANs), diffusion-based models, and large language models (LLMs). Other DGM families, such as normalizing flow models [22], energy-based models [23], and score-based models [24], often integrate with diffusion-based or LLM-based models and are more prevalent in the computer science community. To maintain consistency and avoid misalignment, this review will focus on the aforementioned models, which have already shown promising applications in industrial CM and SHM systems.
Comprehensive scientometric analysis was performed among multiple databases, including Web of Science, Scopus, IEEE Xplore and Google Scholar, to identify relevant publications up to July 2025. Considering that GenAI represents an emerging and rapidly developing research hotspot, numerous studies are available as preprints before being sent to formal peer review. Therefore, the search scope was extended to include repositories such as arXiv, SSRN to capture the most recent trend. The search strategy combined terms related to generative modeling and industrial applications, using keywords such as deep generative model, generative adversarial network, condition monitoring, and structural health monitoring. To ensure completeness, variations and synonyms of these terms (e.g., fault diagnosis, predictive maintenance, and data-driven modeling) were also incorporated. In this way, we provided a broad overview of the evolving scope of generative models applied to industrial health monitoring tasks and systems.
To further elaborate on the co-occurrence among the keywords, we create a map of the cluster of keywords in Fig. 1 using VOSviewer developed by the Center for Science and Technology Studies at Leiden University [25]. The visualization reveals that the generative model exhibits a strong semantic relation with terms such as condition monitoring, fault diagnosis, and data augmentation for vibration signals. However, recent trends indicate a shift toward applications in image processing, synthetic data generation, conditional GANs, and etc. This trend suggests that the deployment of DGMs in industrial CM/SHM remains in its nascent stage, with promising opportunities for future industrial scenarios.
1.3 Motivations
In response to emerging challenges and research trends, this review systematically summarizes DGMs in CM/SHM applications, addressing downstream tasks by identifying faults and detecting possible anomalies. This review provides three key contributions: (i) a thorough summary of state-of-the-art DGMs; (ii) an exploration of their implementations across four technical domains, including data augmentation&imputation, domain adaptation&generalization, multimodal data fusion, and FD&AD tasks in CM/SHM systems; and (iii) a forward-looking perspective on current limitations and research opportunities. This work arrives at a critical moment in the evolution of CM/SHM, where these systems are transitioning from effective signal processing techniques to intelligent, interactive, AI-driven architectures. The emergence of GenAI positions DGMs as transformative tools in this paradigm shift. Building on cross-disciplinary advances from the computer vision to PdM communities, this review establishes a technical framework to address the following questions.
What DGMs have been previously used to support research related to CM/SHM systems?
Previous studies on using DGM in CM/SHM systems mainly focus on DAR, VAEs, and GANs, which are dedicated to data augmentation, domain adaptation and generalization. Among them, i) DAR models are widely used in time series augmentation; ii) VAEs focus on data augmentation through latent space modeling; and iii) GANs enhance unbalanced or missing data by synthesizing limited fault data through adversarial training. Diffusion models and LLMs are recent emerging trends. Diffusion-based models are mainly used for signal denoising and data augmentation, while the application of LLMs is still limited to fault diagnosis and damage detection based on pre-trained large models and fine-tuned on customized datasets.
Why DGMs have great potential to enhance or even replace signal processing-based methods in specific application scenarios?
Unlike signal processing techniques that typically rely on hand-crafted feature extraction methods such as Fourier transform or wavelet transform, DGM can automatically learn hierarchical representations from multisource and multimodal data. This enables DGM to perform end-to-end learning without the need for multi-step processing and feature extraction. However, in resource-constrained environments, or when interpretability and computational efficiency are critical, signal processing techniques may still be preferred.
What are the remaining challenges and future research directions in using DGMs for CM/SHM system?
The high computational cost, difficulty in creating high-quality synthetic data, and interpretability challenges remain the main limiting factors for applying DGM in the field of CM/SHM. Future research can explore physics-based generative networks to improve the consistency and fidelity of generated signals, leverage multimodal data generation techniques, and combine reinforcement learning with LLM to simulate extreme failure scenarios (e.g., catastrophic equipment failures or rare structural failures), thereby improving the model robustness of the entire industrial ecosystem.
The remainder of this paper is organized as follows: Section 2 presents the theoretical background and recent state-of-the-art DGMs. Section 3 reviews the tasks and specific industrial applications. Section 4 discusses existing challenges and future research directions. Section 5 concludes the review.
2 Preliminaries of DGMs
2.1 Theoretical preliminaries
Mathematically, the primary goal of DGMs involves learning representations of intractable probability distributions over , where is typically large and the distribution exhibits complex characteristics. This learning process uses a finite but substantial collection of independent and identically distributed (i.i.d.) samples from , referred to as training data. Unlike conventional statistical inference that seeks explicit probability expressions, the main objective of DGMs is to construct a generator [16]:
where maps data from a tractable distribution supported on into data resembling the target distribution . Formally, let us assume that for each , there exists at least one latent vector satisfying . The transformed distribution thus approximates , which can be illustrated in Fig. 2.
2.2 State-of-the-art DGMs
2.2.1 Deep autoregressive models
Deep autoregressive models belong to a class of DGMs that generate new data samples by iteratively invoking the model based on past observations [26, 27]. This makes them particularly suitable for sequential data modeling tasks such as time series forecasting. Given a dataset with each , the training objective for a DAR model minimizes the negative log-likelihood (NLL) of the data under the model, computed via the chain rule of probability [17]:
where is the conditional probability density for the -th time step given previous steps, parameterized by . denotes the expectation over all samples. In practice, the model learns to maximize the likelihood of each subsequent value in a sequence, given its previous values.
2.2.2 Variational autoencoders
VAEs represent a typical class of DGMs specializing in data compression and reconstruction [17]. Their architecture consists of two networks: an encoder and a decoder. The encoder maps high-dimensional inputs to a lower-dimensional latent space that captures essential data features, while the decoder reconstructs approximations of the original input from these latent representations. This framework achieves effective dimensionality reduction by identifying latent variables that encode the intrinsic structure of data. Formally, a VAE is a probabilistic model that processes an input vector by introducing a latent variable with a predefined prior distribution (e.g., standard Gaussian distribution). Given a dataset with each , the model optimizes the log-likelihood of the data under the generative process:
To tractably maximize this log-likelihood, VAEs optimize the Evidence Lower Bound (ELBO) for each data point , defined as [28]:
where denotes the parameters of the decoder reconstructing data from latent variables. represents the parameters of the encoder approximating the posterior distribution over latent variables. stands for the Kullback–Leibler divergence. By balancing accurate reconstruction with latent space regularization, VAEs efficiently learn compressed representations while maintaining the ability to generate coherent outputs from the latent space.
2.2.3 Generative adversarial networks
GANs [29] are a class of DGMs that employ a unique adversarial training process. A GAN consists of two competing networks: a generator (denoted ) and a discriminator (denoted ). The generator tries to create synthetic samples (e.g., thermal images, vibration acceleration signals) given a random noise, while the discriminator attempts to distinguish between whether the generated samples are real or fake [30]. This adversarial process leads the generator to progressively refine its output samples, ultimately producing realistic high-fidelity outputs. The objective function governing this process is formulated as a minimax optimization problem [31]:
where represents the adversarial loss of GAN, and denotes the data sampled from . is the probabilistic distribution in latent space (typically Gaussian).
2.2.4 Diffusion models
Denoising diffusion probabilistic models (DDPMs) [32], commonly referred to as diffusion models, form the foundation of the diffusion-based GenAI framework. These models represent a class of DGMs that have demonstrated exceptional performance in image generation tasks. The core of the DDPMs is to construct two Markov chains: a forward diffusion process and its corresponding backward denoising. The forward diffusion process incrementally corrupts an initial data sample over discrete steps by introducing additive Gaussian noise, until it is transformed into a pure noise sample . Reversely, the backward denoising, as it indicates, iteratively reverses this transformation to reconstruct the original data distribution from pure noise. Formally, the forward process defines a Markov chain governed by the forward transition probability [33, 32]:
where denotes the time-dependent noise scale at step , is the identity matrix, and represents the Gaussian distribution. The backward denoising denoted as iteratively reconstructs the original data distribution from pure noise . Specifically, the backward process defines another Markov chain governed by [32]:
where the mean and covariance are learned by a neural network (typically using a U-Net architecture), parameterized by . The loss function of training is to maximize the likelihood by maximizing the ELBO, . The ELBO can be written as matching the true denoising distribution with the parameterized denoising model by [34]:
where C is a constant. The forward diffusion and backward denoising processes, and key characteristics of the diffusion model are also summarized in Table 1.
2.2.5 Large language models
LLMs have become a dominant paradigm for addressing various engineering challenges. These models are particularly notable for their ability to process diverse inputs, such as text prompts, questions, or visual images through prompt engineering. Representative LLMs such as the generative pre-trained transformer 4 (GPT-4) 111https://openai.com/index/gpt-4/, Llama 4 222https://www.llama.com/, Claude 333https://claude.ai/, and DeepSeek 444https://www.deepseek.com/, all utilize a Transformer-based architecture pre-trained on massive datasets. This architecture integrates attention mechanisms to capture both local and global dependencies between tokens [35]. Mathematically, LLMs function as autoregressive models that predict the next token conditioned on previous tokens, which can be expressed as the following conditional probability:
where represents the transition probability of the next token given all previous tokens up to the position. denotes the learnable weight matrices, is the hidden state derived from the Transformer by processing the previous input sequence , and softmax refers to the softmax activation function that converts logits into a probability distribution for token selection. This indicates the model will output the token with the highest probability after the softmax function. A comparison of the model architectures and specific characteristics of the DGMs presented is illustrated in Table 1.
Models
Base architectures
Pros
Cons
DAR
Sequential generation given previous inputs
Specialize for time series or flow data [36]
Slow to train and converge
Struggle to model complex, non-autoregressive distributions [37]
VAE
Probabilistic encoder-decoder architecture [38]
Good to learn latent representation
Stable training process
Sensitive to a limited dataset
May suffer from posterior collapse [39]
GAN
Adversarial generator-discriminator framework
Produces high resolution synthetic samples [33]
Prone to training instability [19]
The training loss can be slow to converge [40]
Diffusion model
Forward diffusion and backward denoising process [32]
Stable training via gradually removing noise
High-fidelity sample generation [26]
Large number of denoising steps to cause training complexity [34]
Hard to control over the generation process
LLM
Transformer-based architecture
Pre-trained on large image text paired datasets
Multimodal fusion via attention mechanisms [35]
Large computation resources both in the training and inference stages
Large high-quality training data
∗ The thermal images of the motor illustrated in the table were retrieved from [41].
2.3 DGMs vs Classical generative models
Section 2.2 outlines the basic theory of state-of-the-art DGMs. Although generative models have developed rapidly in recent years, some classic methods remain useful and conceptually significant for CM/SHM. Representative classical generative methods include Gaussian mixture models (GMMs) [42], hidden Markov models (HMMs) [43], and autoregressive time-series models such as ARIMA. These are often referred to as statistical or classical generative models. Classical generative models are parametrically compact: a small number of parameters, low computational cost, strong interpretability, and good performance on small datasets as they can explicitly encode domain-specific structures such as linear dynamics or known statistical dependencies. However, their representational capability is limited, as they are often too shallow to capture the highly nonlinear, high-dimensional, or multimodal patterns present in multi-sensor data.
It is worth noting that the autoregressive models span a wide range of models from these classical autoregressive models through deep autoregressive networks to the latest GPT-style autoregressive large models (i.e., LLMs) [44]. On that basis, we compare classical generative models and DGMs with respect to model capacity, data and computational requirements, inference speed, and interpretability, as shown in Table 2.
Metric Classical generative models Deep generative models (GMM, HMM, ARIMA, etc.) (DAR, VAE, GAN, Diffusion, LLMs, etc.) Training objective Maximum likelihood, Expectation–Maximization, least squares ELBO (VAE), adversarial loss(GAN), denoising score (Diffusion model) Data types Low dimension data, time series High-dimensional complex data: images, audio, multimodal sensor data Model complexity Typical range: [45] Typical range: (VAEs/GANs [46], transformers/LLMs [47]) Evaluation metrics Log-likelihood, RMSE, prediction interval coverage NLL, ELBO, Fréchet inception distance, task-specific scores (F1-Score, AUC, etc.) Inference speed , typically on CPU [48] Normally on GPU; Diffusion/LLM inference: [49] Interpretability Highly interpretable Low/moderate interpretability due to a large amount of model parameters
3 DGMs application in CM/SHM
The previous section introduced the latest DGMs in detail and compared the advantages and disadvantages between classical generative models and DGMs. This section systematically investigates DGMs in different tasks and various industrial systems. Section 3.1 describes the specific tasks performed by DGMs, and Section 3.2 concludes the industrial systems applied in the CM/SHM communities.
3.1 Applications in specific tasks
3.1.1 Performance metrics
Before introducing the DGM application to specific tasks, the quantitative metrics are given in Table 3 to fully evaluate the performance of the DGM. These metrics are widely implemented in various subsequent tasks. For instance, the mean absolute error (MAE) can be used to evaluate the quality of synthesized data compared to real data in data augmentation tasks, as well as in fault diagnosis and anomaly detection tasks. The Pearson correlation coefficient (PCC) is used to evaluate the similarity between real and synthetic data in the augmentation task.
Performance metrics Mathematic expressions Mean absolute error (MAE) Mean absolute percentage error (MAPE) Root mean square error (RMSE) Accuracy (ACC) Correlation coefficient () Root mean square (RMS) Pearson correlation coefficient (PCC) Crest factor Maximum mean discrepancy (MMD) KL Divergence () Structural similarity index measure (SSIM) Signal-to-noise ratio (SNR) Earth mover’s distance (EMD) Classification metrics Accuracy, Precision, Recall, F1-score, confusion matrix
Note: , , , denote the real data, its averaged value, probability distribution and standard deviation, respectively; and represent synthetic data, generated, estimated or imputed data, and its probability distribution.
3.1.2 DGMs for data generation
Data generation serves as an umbrella concept with data augmentation, reconstruction and imputation, which differ mainly in how and why the data are generated. Data augmentation can improve model robustness by generating new samples from existing data, while data reconstruction tries to restore or rebuild signal from partial or compressed data. In contrast, data imputation focuses on filling in missing or corrupted values to complete the dataset. Following the data augmentation methods defined in [50], these techniques can be categorized into three primary groups: the synthetic minority over-sampling technique (SMOTE), data transformation-based augmentation, and generative model-based approaches. SMOTE-based methods specifically address class imbalance by synthetically generating minority class samples (e.g., fault or defect samples in CM/SHM applications). Whereas data transformation-based approaches enhance dataset diversity through operations such as noise injection or image manipulations (e.g., cropping, rotating, flipping, and padding). Fig. 4 is an illustration for both 1D time-series and 2D image augmentation in CM/SHM applications. The third category comprises generative model-based approaches, which in this work refer to DGM-based approaches.
DGM-based data augmentation methods have emerged as indispensable tools to address challenges in scenarios where real-world data is scarce and noisy. These models enable robust supplementation and augmentation of existing datasets by synthesizing high-quality data that mirrors the statistical properties of real data samples, especially when acquiring experimental operation data is often prohibitively costly and labor-intensive [53, 54]. Thus, by adopting DGM-based data augmentation, industrial sectors are expected to mitigate data scarcity and address imbalance issues inherent to industrial settings, thus enriching training datasets for downstream FD&AD models while preserving the integrity of original data distributions [55, 56]. Beyond data augmentation, DGMs also play a critical role in data imputation, a solution to resolve missing data points in industrial systems [12]. Missing data may seriously compromise the diagnostic or detective performance of DGM-based models [57]. The inherent ability of DGMs to learn complex data distributions makes them well-suited for data imputation tasks. Instead of simply filling in missing values with fixed estimates, DGMs can leverage the underlying patterns and correlations present in the data to generate plausible values for missing entries. Table 4 summarizes the existing literature on applying DGMs to data augmentation and imputation tasks. We comprehensively evaluate signal processing, conventional ML-based approaches (e.g., support vector regression (SVR), recurrent neural networks (RNN), long short-term memory (LSTM), gated recurrent unit (GRU)), and DGM-based methods using the performance metrics outlined in Table 4. SMOTE-based approaches are classified as conventional ML, as SMOTE generates synthetic data through linear interpolation between minority class samples and their k-nearest neighbors (KNN) [58].
Existing studies explicitly focused on signal processing, conventional ML, and DGM are listed in Refs. Works comparing the performance improvements across these approaches using evaluation metrics in Table 4 are detailed in Quantitative comparison. From the table, it is found: (i) most works focus exclusively on data augmentation or imputation without cross-comparative analysis; (ii) in quantitative comparisons, DDPM and GAN-based models consistently outperform conventional ML methods (e.g., LSTM, SMOTE) across metrics such as MMD and EMD; (iii) cross comparisons between signal processing methods, conventional DNNs, and DGMs remain scarce, largely because signal processing techniques are primarily employed as preprocessing stages.
Moreover, diffusion-based models are typically large-scale architectures containing billions of parameters. Recent studies have examined the relationship between the training time and data generation quality in diffusion-based models [59, 60, 61]. As discussed in these works, the number of diffusion steps is a key factor determining the final data generation quality, which typically ranges from 500 to 1000 steps [61]. This observation has motivated the development of denoising diffusion implicit models (DDIMs) [34], which aim to accelerate the sampling process while maintaining generation fidelity. In practice, the training time for a typical diffusion model often exceeds 10 hours even when using commercial-grade GPUs.
Signal processing Conventional ML/DNNs DGMs Data transformation-based/Least-squares SMOTE/CNN/RNN/LSTM DAR/VAE/GAN/Diffusion Data Augmentation Refs: Data transformation-based [62, 63, 64, 65] Refs: SMOTE-based [66, 67, 68] Refs: DAR-based [69, 70, 71, 72], VAEs [73, 74], GANs [75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90], Diffusion-based [91, 92, 93, 60, 94, 95, 96, 97, 98, 99, 100, 101, 102] Quantitative comparison: – [103] DDPM-based method surpasses GAN-based and autoencoder model in augmentation task with metric: MMD . – [104] Different GAN-based variants for wavelet transformed image augmentation: MMD, EMD, , SSIM. – [105] Comparison of LSTM, VAE, GAN-based data augmentation: Improved diagnosis performance ACC (confusion matrix). – [106] SMOTE and GAN-based data augmentation performance comparison: classification accuracy (F1-score, confusion matrix). Data Imputation Refs: Least-squares [107], Bayesian [108] Refs: KNN [109], Principal component analysis (PCA) [110], RNN [111] Refs: DAR-based [112, 113], VAEs [114], GANs [115, 116, 117, 118], Diffusion-based [119] Quantitative comparison: – [120] The proposed LSTM-based methods outperform random forest (RF), light gradient boosting machine, and multiple linear regression for dams missing data imputation task: , MAE , RMSE . – [121] The proposed Bayesian matrix learning approach outperforms PCA in SHM data imputation: higher ACC (0.95). – [122] Comparison among KNN, RF, matrix factorization, RNN, their proposed model: MAE (7.77), RMSE (10.92). – [123] Dam missing value imputation among GRU, LSTM, and their proposed model: MAE (0.593), MAPE (0.076), RMSE (0.731). – [124] GAN surpasses SVR, radial basis function network, and RNN in data imputation: lower MSE (0.0237), higher (0.9946). – [125] The proposed GAN-based method surpasses LSTM and SVR in bridge data imputation: ACC (0.9), lower execution time. – [126] The method outperforms the VAE and other GAN variants with: lower RMSE (), MAE () values, and higher . Domain Adaptation Refs: Not applicable Refs: Gaussian mixture model [127], CNN-based models [128] Refs: GANs [129, 130, 131, 132, 133, 134, 135, 136, 87], Diffusion-based [137] Quantitative comparison: – [138] Comparison among different DA methods on four SHM scenarios: ACC , F1-score (1.000). – [139] Improved diagnostic performance of proposed dynamic collaborative adversarial domain adaptation network in unsupervised DA task: recall , precision , accuracy , F1-score (classification metrics). – [140] Multi-source DA comparison among diffusion and DNN models: improved classification performance (confusion matrix, accuracy). Domain Generalization Refs: Not applicable Refs: Mete-learning techniques [141] Refs: GANs [90], Diffusion-based model [142] Quantitative comparison: – [143] The proposed adversarial domain generalization framework that outperforms LSTM, CNN-based, temporal CNN, and multiscale CNN with evaluation metrics on unknown target domain: RMSE (0.1376), MAE (0.1117). – [144] Meta-learning-based DG model outperforms domain adversarial neural network, correlation alignment, meta-learning DG, aggregation with improved metrics: accuracy (), enhanced classification performance with confusion matrix. – [145] The proposed similarity-based meta-representation learning method for DG task outperforms several baseline methods including neural network, hierarchical neural network, model-agnostic meta-learning (MAML), and hierarchical MAML with improved metric: ACC .
3.1.3 DGMs for domain adaptation and generalization
Transfer learning (TL) aims to leverage knowledge from a pre-trained model on a source task to improve performance on a different yet related target domain or task [146, 147]. Within this framework, domain adaptation (DA) and domain generalization (DG) emerge as two pivotal subfields. DA focuses on minimizing data distribution discrepancies between the source and target domains [148, 149], while DG seeks to transfer knowledge from multiple source domains to an unseen target domain without utilizing any data from the target domain. Hence, the main difference between DA and DG lies in data accessibility during training: DA can utilize target domain data during training (even when such data are scarce or limited), while DG relies solely on source domain knowledge. This makes DG particularly suitable for evaluating the robustness of the model in completely unseen working scenarios where the target domain data remain inaccessible [82]. Fig. 5 illustrates the main difference between DA and DG concepts.
Following the framework established in the previous subsection, Table 4 outlines the applications of DGMs for DA and DG within the TL framework. Apparently, signal processing methods are not applicable to either DA or DG, as they belong to transfer learning, which is indeed one category of ML methods. Nevertheless, recent work mainly focuses on GAN-based models for DA, including: (i) undamaged-to-damaged data transition under scarce damage data [129, 130, 131]; (ii) adaptation between simulation and experimental data [135, 132, 139], and (iii) adaptation among different rotating machine types, speeds, or operating conditions [133, 134, 87]).
A smaller proportion of studies [90, 142, 145, 144, 143] address DG under unknown or unseen working conditions. Meta-learning [145, 144] stands out as a complementary method to DG by training a model to adapt quickly to new tasks or environments by leveraging knowledge gained from multiple related tasks. Since DG is practical in industrial settings, where target domain data is unavailable, there is an emerging trend toward the use of diffusion model-based DG for CM/SHM systems [103]. It is noted that certain citations span multiple application categories. For instance, [90, 87, 103] proposed methodologies that combine data augmentation with domain generalization to handle unknown target domains. As shown in Table 4, these approaches inherently span both categories since they both involve augmenting scarce data while achieving data generalization to unseen domains.
3.1.4 DGMs for multimodal data fusion
In the context of multimodal data fusion in condition monitoring, the application of DGMs remains in its early stages. Notably, diffusion-based and LLM-based DGMs exhibit unique promise largely because of their Transformer-based architectures, which inherently enable the dynamic fusion of heterogeneous data types, including text, visual images, and audio inputs (known as multimodal data), into a unified latent space. The attention mechanism [35] within these Transformer frameworks serves as a foundation for cross-modal fusion. By dynamically computing attention scores across modalities, this mechanism enables context-aware weighting of input features, allowing the model to prioritize statistically salient or semantically correlated signals during fusion. These capabilities are particularly important for diffusion-based and LLM-based DGMs, as the attention mechanism not only enhances representational power but also ensures the coherence of model outputs with the joint distribution of multimodal inputs.
Fig. 6(a) provides an illustrative scheme of the multi-head attention mechanism in the Transformer network, which is a fundamental component in both Diffusion-based (e.g., Stable Diffusion [150]) and LLM-based models. Fig. 6(b) illustrates the implementation of cross-modal fusion in LLMs, where paired image and text inputs are jointly processed. In this framework, the image (or visual) inputs are sent to a feature extraction module via an image encoder to obtain latent representations, which are subsequently projected into key () and query () matrices. Meanwhile, the text caption is encoded by a text encoder (e.g., CLIP text encoder [151]) into a value () matrix, as shown in Fig. 6(a). In this way, the LLM computes attention scores by enabling the dynamic fusion of image and text modalities. Fig. 6(c) depicts the main flow of multimodal data fusion based on DGM. Various data formats, such as discrete data, continuous time series, or images, can be obtained from multimodal heterogeneous sensors (such as vibration accelerometers, stereo cameras, and acoustic emission sensors). The DGMs can operate on and dynamically fuse this information to facilitate downstream FD&AD tasks. Table 5 concludes the DGM application for multimodal data fusion.
In Table 5, we categorize diffusion-based and LLM-based models derived from Transformers as DGMs, while other Transformer-based networks are classified as DNNs. Within multimodal data fusion, diffusion and LLM-based models constitute a significant proportion of DGMs. Conversely, works using Kalman filtering for multimodal fusion are classified as signal processing methods. Although Kalman filtering is a popular and effective approach for real-time multimodal fusion in monitoring scenarios, it struggles to handle large volumes of heterogeneous data streams. This limitation indicates the potential for applying diffusion and LLM-based models to future multimodal data fusion techniques.
Signal processing Conventional ML/DNNs DGMs Kalman filtering/Least-squares CNN/RNN/LSTM/Transformer Diffusion/LLM Multimodal Data Fusion Refs: Kalman filtering [152, 153, 154] Refs: LSTM-based [155], Transformer [156, 157] Refs: Diffusion-based [158], LLM-based [159, 14, 160, 161, 162] Quantitative comparison: – [163] A multi-scale Transformer network was proposed with acceleration feature fusion achieving improved performance comparing to CNNs and sigle-scale Transformer network: Accuracy (96.12%), confusion matrix. – [164] The proposed Swin-Transformer-based model achieves highest diagnostic accuracy compared to ResNet18, EfficienceV2, and Vgg16 networks by fusing infrared thermal and visible light images: Accuracy , confusion matrix. Fault Diagnosis &Anomaly Detection Refs: Not applicable Refs: CNN-based [165, 166, 167], Transformer [168, 169, 156, 157, 170, 171] Refs: Diffusion-based, LLM-based [14, 160, 161, 162] Quantitative comparison: – [156] A two-stage encoder Transformer was proposed with 50% improvement in accuracy comparing to extreme gradient boosting. – [157] The temporal fusion Transformer network achieves higher accuracy compared to the autoregressive model: MAE, MSE, MAPE. – [172] The proposed model outperforms ShuffleNet, ResNet, DenseNet models with improved: average accuracy ().
3.1.5 DGMs for fault diagnosis and anomaly detection
For downstream FD&AD tasks, it needs to be noted that the CM/SHM process collects sensor data from different positions and channels with varying types and sampling rates. FD and AD were carried out based on the collected data. LLM-based diagnosis and prognostic methods have recently been developed, thanks to the development of game-changing products, such as ChatGPT, which are developed on Transformer-based architecture with the ability to create text, images, and even videos.
The basic idea behind applying LLM-based models in FD&AD tasks is that the Transformer architecture in LLMs integrates a large number of model parameters that are pre-trained on large text-image datasets. Then the pre-trained LLMs can be adapted to bearing or structural FD&AD tasks by fine-tuning the customized dataset. As shown in Fig. 7, the authors conducted fault diagnosis by using Low-Rank Adaptation (LoRA) fine-tuned GPT-2 model based on bearing vibration datasets. In Table 5, we summarize the LLM-based FD&AD tasks presented for the wind turbine system [14], and the bearing/gearbox [160, 161, 162]. Most of the publications for LLM-based FD&AD tasks focused on bearing/gearbox and were also based on LoRA fine-tuning techniques. We also detailed the different types of fine-tuning techniques in the following Section 4.1.1.
It is noted that the fine-tuning efficiency and model size of LLMs largely determine their final performance in FD&AD tasks. Zheng et al. [173] compared the diagnostic accuracy of Llama-2 models of different sizes (7B vs 13B), and reported that the 7B model can achieve satisfactory results, while larger models are generally preferred for tasks requiring higher precision or involving more complex datasets. Additional studies discussing the impact of fine-tuning strategies and model scale on LLM performance for FD&AD can be found in Refs. [174, 175, 176, 177].
3.2 Applications in industrial systems
In the previous subsection we summarize the main tasks performed by DGMs. To illustrate how these tasks transfer into practical use, this section describes DGM applications across subsystems and components in four major industrial sectors: rotating machinery, aircraft structures, wind turbines, and civil infrastructure. Fig. 8 maps base DGM families and their variants to the tasks they perform and the subsystems they enable in CM/SHM. From the figure, it can be observed that GAN and its variants dominate in data generation task and most of tasks are applied to rotating machinery and civil infrastructure sectors. The diffusion-based models and LLMs remain in the early stages of industrial practice.
3.2.1 Rotating machinery
Rotating machinery is a major and critical component of many mechanical systems, serving primarily to generate, transmit or regulate rotational motion and torque [178]. A representative example is the gearbox, a mechanical device designed to adjust the rotational speed and output power to the desired level. Inside the gearbox, gears transmit torque through precise meshing, while bearings support power transmission by reducing friction and stabilizing the rotating shafts. These components, however, are highly susceptible to unexpected failures: bearings are likely to experience inner race, outer race, or rolling element faults [78], whereas gears are prone to surface fatigue (pitting, case crushing, and spalling) and even tooth breakage [179]. Therefore, continuous monitoring is essential to detect problems and predict potential failures.
As shown in Table 6, DGMs have been extensively applied to vibration acceleration signals, supporting a range of tasks discussed in Section 3.1. At the component or subsystem level, gearboxes and bearings are the most representative components in rotating machinery. GAN and diffusion-based models are commonly employed for data generation, domain adaptation, and generalization tasks. Recently, LLM-based approaches for multimodal fusion and fault diagnosis have emerged as a growing trend. The text prompting functionality enabled by LLMs allows the recording and interpretation of metadata, such as operational dialogues and state changes, which other DGMs cannot process, thereby offering new opportunities for knowledge-enhanced condition monitoring.
3.2.2 Aircraft structures
Aircraft SHM concentrates on ensuring structural integrity of load-bearing elements such as wings, fuselage panels, and landing gear structures. The composite wings and fuselage of commercial aircraft are subject to extreme environmental conditions from sub-zero temperatures at cruise altitudes to intense solar heating and thermal cycling on the ground. Continuous monitoring and maintenance in such environment are exceedingly difficult. Typical sensing approaches primarily rely on NDT techniques, including ultrasonic sensing and visual inspection, to detect potential damage such as delamination or fiber breakage. DGMs assist in aircraft SHM by generating realistic signals and features for scarce experimental dataset as shown in Table 6, enabling cross-modal synthesis, and producing conditioned damage progression data for downstream diagnostic and prognostic purposes [61].
It should be noted that engine health monitoring is an extensive topic treated under energy-based studies and will not be discussed in this work.
3.2.3 Wind turbines
Wind turbines represent a crucial source of renewable energy. In their operational state monitoring, Supervisory Control and Data Acquisition (SCADA) systems continuously record a wide range of parameters in real time, including wind speed, gear oil temperature, lubricant pressure, generator output power, and rotor speed. At the system level, SCADA data consists of comprehensive state monitoring parameters, aligning closely with the CM community. It should be noted that the rotating components such as the gearbox in wind turbines are classified as DGM-based applications in rotating machinery as discussed in Section 3.2.1.
At the component level, blade inspection constitutes a key focus area within the SHM community. The turbine blades are constantly subjected to environmental stress from wind, rain, sand, and other particles, leaving them extremely susceptible to wear and tear [180]. Due to these harsh conditions, detecting blade damage using vibration or vision-based inspection becomes a critical task in wind turbine health monitoring.
3.2.4 Civil infrastructures
Civil structure monitoring addresses the safety, serviceability, and longevity of large-scale infrastructure systems such as bridges, tunnels, structures, dams and etc. At the system level, SHM emphasizes long-term sensing and decision-making to manage aging, extreme loadings, and changing operational demands.
At the component level, subsystems of interest include bridges, tunnels, dams, and buildings. Common modalities span point sensors (strain gauges, accelerometers), distributed fiber optic sensing systems, NDT techniques (guided waves, acoustic emission), and remote imaging (UAV photography, LiDAR). DGMs complement these capabilities by enabling multimodal fusion, robust data imputation, synthetic generation of damage scenarios (such as cracks, corrosion), conditional/physics-aware simulation for digital twin calibration, domain adaptation across sites and sensor layouts, virtual sensing to estimate unmeasured fields, etc. Although UAV-based inspections are increasingly important for dams and large-span structures, vibration and strain-based data generation tasks remain prevalent across civil applications. The mapped DGM tasks and component-level applications are summarized in Table 6.
Industrial sectors Subsystems DGMs Data types Specific tasks Rotating machinery Gearboxes, Bearings VAE [181, 182, 183] Vibration; Acoustics Fault diagnosis GAN [78, 79, 80, 85, 81, 82, 83, 84, 86, 90, 87, 88, 89, 184, 185] Vibration; Acoustics; Thermal images Data generation GAN [132, 139, 143, 90, 133, 134, 87, 186] Vibration; Acoustics Domain adaptation and generalization Diffusion [95, 103, 96, 97, 98, 99, 100, 101, 187] Vibration; Acoustics Data generation Diffusion [140, 103] Vibration Domain adaptation and generalization LLM [160, 161, 162] Vibration; Text prompt Multimodal fusion LLM [160, 161, 162] Vibration; Text prompt Fault diagnosis, anomaly detection Aircrafts Blades GAN [188] Loading data Data generation Landing gears GAN [189] Landing gear data Data generation Wings, skins GAN [190] Visual images Data generation Diffusion [191, 192] Ultrasonic signals Data generation Wind turbines VAE [193, 194] SCADA GAN [115, 195] SCADA data Data generation Diffusion [196, 158] Blades VAE [197] Vibration Anomaly detection Diffusion [198] Visual images Anomaly detection Civil structures Bridges DAR [69] Strain Data generation VAE [199] Vibration Anomaly detection GAN [116, 124, 76, 117] Vibration; Strain; Visual images Data generation GAN [135] Vibration Domain adaptation and generalization Diffusion [91, 60, 200] Vibration Data generation Tunnels GAN [75] Vibration Data generation Buildings GAN [201] Visual images Data generation GAN [129, 130, 131] Vibration Domain adaptation and generalization Dams GAN [202] Visual images Anomaly detection
4 Challenges and future prospects
This section summarizes the current limitations of DGMs in CM/SHM into three main aspects, as described in Section 4.1: the need for efficient fine-tuning strategies, interpretable and reliable models, and mitigating the computational inefficiency of edge deployment. Corresponding future prospects will be discussed in Section 4.2.
4.1 Challenges and limitations
4.1.1 Fine-tuning techniques
DGMs are generally trained on large datasets during the initial pre-training stage. To adapt these pre-trained DGMs to specific tasks or domains, fine-tuning techniques are usually adopted by further training on a smaller, custom dataset. By leveraging knowledge from pre-training while learning nuances of the target domain, fine-tuning bridges the performance gap between the source and target domains. This also indicates that fine-tuning is among the most widely used strategies for applying DGMs in transfer learning and domain adaptation across diverse applications. Fine-tuning techniques for DGMs consist primarily of two branches: 1) Full fine-tuning, and 2) parameter-efficient fine-tuning (PEFT) [203]. Full fine-tuning updates all parameters of the pre-trained model while preserving its original architecture, and the size of the model after fine-tuning remains the same as the pre-trained model. However, this method requires a large storage space for DGMs and is also not achievable for deploying the fine-tuned models for downstream applications. While PEFT methods selectively update a small subset of parameters, drastically reducing computational and storage overheads [204]. PEFT methods are further categorized into:
Prefix/Prompt-Tuning [205]: Additional trainable prefix tokens added to the input or hidden layer (these prefixes are continuous pseudo tokens that do not correspond to real tokens), and only these prefix parameters are trained.
Adapter-Tuning [206]: Insert smaller neural network layers or modules into each layer of the pre-trained model. These newly inserted neural modules are called adapters. When fine-tuning downstream tasks, only these adapter parameters are trained.
Low-Rank Adaptation [207]: Learning a low-rank matrix with small parameters to approximate the parameter update of the model weight matrix, only the low-rank matrix parameters are optimized during training.
As illustrated in Fig. 9, LoRA operates by decomposing model updates into low-rank matrices, enabling efficient adaptation without altering the original parameter count.
In industrial applications, efficient fine-tuning enables pre-trained DGMs to generate domain-specific sensor data (such as vibration accelerations, infrared images, and acoustic emission signals) while maintaining their foundational understanding of complex patterns [61]. For instance, a DGM pre-trained on diverse mechanical fault datasets can be fine-tuned with LoRA or adapters based on the Hugging Face PEFT package [203] using only a small number of labeled data from a particular wind turbine gearbox. This enables the model to detect subtle anomalies on-site without requiring a complete retraining of the model from scratch, which is a promising direction for real-time condition monitoring and possible deployment in industry.
4.1.2 Explanability, interpretability, and trustworthiness
Explainable models pose another obstacle in deploying DGMs for industrial generative tasks. Since DGMs combine generative modeling with DNNs, which are often characterized by the notion of black-box models, it is important to explain the generated results and interpret internal DGM mechanisms, which are collectively termed as Explainable Generative AI (GenXAI) [208]. Note that explainability and interpretability represent distinct but related concepts, where the former tries to explain and illustrate model decisions or outputs in human-understandable terms [209], whereas the latter emphasizes understanding the internal processes governing model behavior [210]. In contrast to traditional DNN models, the scale of modern DGMs, particularly LLMs, with billions of parameters and vast training datasets introduces huge challenges for GenXAI research.
To date, different explainable skills have been developed in the realm of GenXAI, which can be primarily categorized into two types of methods: local explanation (seeks to explain the result) and global/mechanistic explanation (interprets its internal mechanism) [211]. Local approaches often adopt techniques from traditional DNN explainability, such as feature attribution methods (in this case, SHapley Additive exPlanations (SHAP) [212]). More recent local explanation techniques use LLMs themselves to yield explanations, e.g., through post-hoc natural language explanations, asking an LLM to build explanations into its generation process, or through data grounding (e.g., Retrieval-Augmented Generation (RAG) [213]). As an alternative, the second direction is for global or mechanistic explanation, which also means to interpret the entire model. Similar techniques have also been developed, such as attribution of DGM internals, interpretation of model internals, algorithmic understanding, and data influence [214]. To conclude the aforementioned content, we elaborate on the difference between local and global explanation as shown in Fig. 10 from four well-built facets: attribution, natural language, decomposing reasoning, and data grounding.
Trustworthiness remains a critical concern, particularly for LLMs. While these models are good at generating accurate predictions by leveraging statistical patterns and semantic patterns from training data, their reliability is inherently tied to the limitations of DNN models [215]. In LLMs, the models may get a good handle on grammar, syntax, and semantic relationships in language usage, but they may not necessarily map the words to real-world entities as humans do, which is crucial for ensuring the credibility and trustworthiness of the generated results. Instead, current LLMs can be prone to hallucinations, which result in the generation of content that deviates from facts or user input [216]. This discrepancy could pose a serious obstacle to practical implementation and cast doubts about the reliability of their diagnostic results.
4.1.3 Quantization and deployment
Deployment involves transferring the algorithm from a development environment into a real-world operational setting where it can start monitoring equipment. Model quantization is an important step before deploying large DGMs on edge devices (e.g., mobile devices on-site) to enable real-time monitoring. The key issue is to eliminate the heavy computational burden caused by the large number of parameters of DGMs, in particular LLMs [217]. At the very beginning, the fine-tuned DGMs need to be quantized so that 16-bit floating-point models are converted into a compact low-bit (4-bit or 8-bit) fixed-point format. This transformation needs to achieve a balance between the efficiency of fixed-point computation and the benefits of reduced memory usage. Other notable techniques include model pruning and knowledge distillation, both of which aim to reduce network size while preserving as much of the original model performance as possible [218]. The main steps involved in this process include: 1) Efficient fine-tuning of DGMs on the custom dataset to adapt to specific application scenarios; 2) Performing post-training quantization (PTQ) after fine-tuning, and the DGMs are quantized to lower accuracy, that is, a 4/8-bit fixed-point model; 3) Suitable model pruning to deploy the fine-tuned model on edge devices. Quantized LoRA (QLoRA) [219] is an efficient quantization method for LLM-based LoRA fine-tuning, as shown in Fig. 11. Following these steps, the pre-trained large DGMs can be applied to commercial-grade chips so that their generative ability can be used in real-world condition monitoring scenarios.
4.2 Future outlooks
4.2.1 Zero-shot/few-shot/multimodal data generation
DGMs, particularly LLMs, are revolutionizing zero-shot, few-shot, and multimodal data generation paradigms by enabling robust and diverse generalization in data-scarce scenarios. Zero-shot learning [146] focuses on diagnosing or predicting unseen samples, which seamlessly refers to domain generalization in Section 3.1.3. Few-shot learning, on the other hand, aims to achieve high performance with only a small number of training samples, often employing meta-learning techniques [220, 221], which is a broader approach where models learn to learn, enabling them to adapt to new tasks quickly with minimal training data. It can be foreseen that DGMs can be pre-trained on large text-sensor data pairs and then can be exploited for zero-shot or few-shot reasoning for either generative tasks or diagnostic tasks. Multimodal generation will also be greatly enhanced with DGMs, especially LLMs [31]. It is interesting to note that in Section 3.1.4, the multimodal data fusion aims for multimodal input data, which herein DGMs similarly enable effective generation of multimodal output data. In industrial systems, DGMs can generate diverse data streams, such as vibration spectra, thermal imaging, acoustic emissions, and even operational logs, which collectively indicate equipment health.
4.2.2 Physics-informed DGMs
Physics-informed neural networks (PINNs) [222] integrated with DGMs represent a promising direction for next-generation condition monitoring systems. By embedding the governing equations of the monitored system as regularization terms or loss functions, PINNs regulate the generated data to align with in-situ physical dynamics. In particular, combining physics-aware principles or rules with LLMs creates new opportunities to simulate industrial phenomena, such as variable loading conditions, thermal gradients, and mechanical stresses. To date, no published research has explored the application of physics-informed LLMs for industrial condition monitoring. Future work might focus on hybrid architectures in which PINNs guide LLMs to generate physically consistent degradation trajectories (e.g., bearing wear or crack propagation) for robust predictive maintenance. Additionally, enhancing physics-informed latent spaces by directly linking generative features to measurable physical parameters (e.g., strain, temperature) would strengthen the interpretability and reliability of generated data.
4.2.3 Reinforcement learning with DGMs
Reinforcement learning (RL) has proved highly effective in computer vision for autonomous driving and robotics. The application of RL for real-time monitoring is still in its initial stages. RL is known for its autonomous and adaptive learning features, and has been instrumental in the training and optimization of large-scale models as well as in solving goal-directed decision-making problems [223]. However, conventional RL struggles with sparse rewards, high-dimensional state spaces, and safety-critical constraints [224], but DGMs can address these challenges by modeling system dynamics, simulating environments, and generating synthetic failure samples. For instance, LLMs can learn latent representations of equipment degradation patterns from historical sensor data, allowing RL agents to train in simulated environments that capture rare fault modes or extreme operational conditions. This reduces reliance on real-world data while improving robustness to unseen anomalies. By integrating LLMs with RL agents, systems can achieve policy optimization through the way in which LLMs predict future states under different maintenance actions, enabling agents to explore strategies safely and efficiently. For partial observability issues, DGMs reconstruct missing sensor signals or infer hidden system states, enhancing RL decision-making under uncertainty. Future work may focus on hierarchical RL frameworks to coordinate long-term maintenance with short-term anomaly detection.
5 Conclusions
This work systematically reviews recent advances and emerging trends in deep generative models (DGMs) for industrial condition and structural health monitoring systems. We examine state-of-the-art DGMs, including deep autoregressive models, variational autoencoders, generative adversarial networks (GANs), diffusion-based models, and large language models (LLMs), and categorize their applications into four critical areas: (i) data generation, (ii) domain adaptation and generalization, (iii) multimodal data fusion, and (iv) fault diagnosis and anomaly detection (FD&AD) tasks. The review addresses key challenges across diverse industrial domains, such as rotating machinery, aircraft structures, wind turbines, and civil infrastructure. Challenges still exist, including the need for efficient fine-tuning strategies, enhanced model interpretability, and optimization techniques such as quantization for edge device implementation. In short, the following main conclusions can be drawn from this review.
GAN-based approaches dominate in data generation task, whereas diffusion-based models and LLMs remain in the early stages of industrial practice.
DGMs for domain generalization need to be further explored, especially under unseen scenarios.
Multimodal data fusion in LLM can be sensibly integrated with real-time monitoring data in the condition monitoring process.
For FD&AD tasks, the LLM can fuse different data streams such as vibration, acoustics, and even operational logging texts, but effective fine-tuning such as LoRA still needs to be in place.
For future perspectives, DGMs are still in their infancy for industrial condition and structural health monitoring; this review provides academic and industrial insights for both research and practical implementation.